OpenStudy (anonymous):

C2h6O(g) + xO2(g) arrow 2CO2(g) + 3H20(l) What i the value of X when the equation is correctly balanced?

OpenStudy (abb0t):

It sort of looks like a combustion reaction, in which case you're missing some stuff. Can you fix it a little so I can give you a better explanation of how to go about this?

OpenStudy (anonymous):

C2h6O(g) + xO2(g) arrow 2CO2(g) + 3H20(l)

OpenStudy (abb0t):

Yes. Combustion is what I thought. Well, to balance this, it might help to make a chart which shows how many of each atom you have on the proiduct and reactant side. Remember, to have an equation balanced, what you need to do is have the same amount of atoms on both sides. Also, note that it follows the distributive property from math. If you add a constant coefficient, like this: \(2^*A_2B_7\) what that means if you added that 2 i put the star on, instead of 2A's, you now have 4 A's and 14 B's. I hope that makes sense.

OpenStudy (abb0t):



OpenStudy (abb0t):

Remember, what you do to ONE side, you must also do to the other to get BOTH numbers on both sides the same!. Notice how you don't have the same number of oxygens on both sides? What two numbers can you multiply to give you both the same number on BOTH sides. Give it a try, and if you still need help, let me know

OpenStudy (anonymous):

Carbon dan Hydrogen have balanced for oxygen at the reactant 1 + 2x at the product 2 x 2 + 3 = 7 hence, 1 + 2x = 7 2x = 7 -1 = 6 x = 6/2 = 3